# JWST Reveals Varied Origins Between Jupiter’s Irregular Satellites

### v1

@@ 3.1 Continua Definitions
The processed, thermally-corrected reflectance spectrum of each target is shown in Figure 3 (Left). Proper analysis of absorption features requires clear definitions of the spectral continuum. To aid reader comparisons, we define a “global” continuum, used in constructing Figure 3 (Right). This continuum was fit per-object with a low order (2-4) polynomial on either side of the identified 3 micron complex. Dividing the spectrum by this polynomial transforms the spectra from relative reflectance to continuum-divided reflectances. In continuum-divided space, spectral features are identified via strong divergence from unity.
@@
All other bands are quantified by fitting a local linear continuum on either side of the feature, then fitting a Gaussian to identify the center and depth from the local continuum. Due to its asymmetry, we quantify Himalia’s 2.7 feature uniquely. Its center is defined by fitting a Gaussian near its sharp minimum, and its depth is defined relative to the linear continuum of the overall 3 band complex.
@@ 3.2 Spectral Groups
We find that the irregular satellites display a variety of spectral characteristics, with some groupings clustered within individual families and some that appear in multiple families. We describe four spectral categories to illustrate these complex relationships. Band parameters for all features observed on individual satellites are given in Table 2. The first group is represented by Sinope, Carme, and Themisto, which are all Trojan-like objects all in different orbital families. Ananke comprises the second group, displaying a deeper, rounded 3.0 micron band. The third group includes Pasiphae, a retrograde satellite, and Lysithea, a prograde satellite, which have similar 3.0 micron features with several weak local minima between 2.7-3.3 . Finally, Himalia and Elara, both in the prograde Himalia family, show a unique band complex that includes a sharp 2.7 micron band transitioning into a broad, round feature from 2.8-3.5 microns.
Sinope and Carme have similarly steeply increasing spectra from 0.7-2.5 microns, and shallow absorption features near 3.0 and 3.4 , akin to the “red” class of Jupiter Trojans (Emery et al., 2011; Wong et al., 2024). Notably, Themisto, a small () individual prograde satellite interior to the Himalia family, groups closely with these more weakly bound retrograde satellites. The spectral slopes of these objects, when overlapping, are consistent with previous measurements at VNIR wavelengths by Sharkey et al. (2023). Carme has a slightly deeper band at 3.0 microns than Sinope and Themisto (depths of vs. ), but they all have similar 3.4 depths ().
@@
Attribution | Organics | Organics | Organics | Organics | Organics | - | Organics | - | |
@@
Ananke is unique within our sample with a deep, rounded 3.0 complex and the most neutral (least-red) slope we observe. Its 3 band depth () is similar to the redder Pasiphae and Lysithea, though their band centers and shapes differ. A narrow feature at 2.86 is overprinted within the rounded 3 complex. Beyond 3.6 , Ananke’s spectrum is neutral and featureless. Our observations of Ananke differ from the ground-based observations of Sharkey et al. (2023) from 0.7-1.6 microns, which showed a slope that is approximately 10% redder. A similar discrepancy was noted by Wong et al. (2024) when comparing their JWST observations of the Trojan Polymele to Sharkey et al. (2019)’s ground-based observations with the same methodology and setup. Given that both Polymele and Ananke have similar brightnesses V 19, it is possible that those short wavelength slope discrepencies are indicative of systematic errors in the ground-based observations of these very faint objects.
Pasiphae (Pasiphae family) groups closely with Lysithea (Himalia family) in terms of spectral slope and absorption features. They trace closely together shortward of . Pasiphae displays a steeper red slope from 3.5 to 5.2 , while Lysithea rolls over to neutral (flat) reflectance, similar to Ananke. Pasiphae and Lysithea display deep absorption features () near 3.0 that are distinct from any other satellites.
A comparison of Pasiphae’s and Lysithea’s continuum-divided bands are shown in figure 4, highlighting the similar absorption shapes between them. Both Pasiphae and Lysithea have additional faint features of 3% overprinted near 3.01 . Lysithea has an additional feature near 3.9 , which is similar to features caused by carbonates. However, carbonates also display absorptions near 3.4 which we do not detect, confusing a clear identification. Lysithea also shows a narrow feature at 2.63 whose origin is less clear, but is consistent with features seen in some carbonates (Bishop et al., 2021).
@@
We also find a feature at 4.25 on Himalia and Elara. This feature is similar to one discovered on the Jupiter Trojan Eurybates by Wong et al. (2024). Figure 5 compares our observations of Himalia and Elara with the the Wong et al. (2024) observations of Eurybates, as processed via our reduction pipeline. We reprocess this spectrum only for methodological consistency between these datasets but do not note any differences between reductions. This feature is similar to those caused by solid-state CO2, which is noteworthy since CO2 is volatile and unlikely to be stable as a crystalline ice on the relatively warm airless surfaces of Jovian satellites or Trojans (Wong et al., 2024). Since CO2 is implausible as pure surface ice, it is likely complexed with more refractory components, similar to the Galilean moons (e.g., McCord et al., 1998; Villanueva et al., 2023; Cartwright et al., 2024). The discoveries of the 4.25 feature on Eurybates, Himalia, and Elara demonstrates that complexed CO2 is present on a variety of objects captured by Jupiter.
The 3.0 absorber(s) for Carme, Sinope, Themisto, Pasiphae, and Lysithea are consistent with a variety of OH-bearing species, which we discuss in Section 8 in the context of meteorite comparisons. The absorption at 3.4 on Elara, Pasiphae, Ananke, Themisto, Sinope, and Carme is consistent with complex organics. All three objects in the Himalia family (Himalia, Elara, Lysithea) display a feature near 3.9 . On Elara, the 3.4 and 3.9 features may be consistent with carbonates, but with an unclear relationship to the 3.9 absorbers on Himalia and Lysithea.
@@ 4 Ammoniated Phyllosilicates in the Himalia Family
Some meteorite materials with significant amounts of adsorbed terrestrial water can reproduce the overall width of Himalia’s band complex. However, we note that experimental studies of both meteorites and terrestrial phyllosilicates find that adsorbed water, and the spectral feature it causes, is easily removed when samples are treated under proper vacuum and desiccation procedures designed to mimic asteroid surface conditions (Beck et al., 2010; Takir et al., 2013, 2019). We therefore disfavor fits to Himalia that use phyllosilicates containing adsorbed water.
@@
Himalia’s combination of absorptions at 2.7 and 3.05 is reminiscent of Ceres (Figure 6), where the 3.05 features is associated with the presence of NH4 in the clay mineral structure (e.g., King et al., 1992; Rivkin et al., 2006; Ammannito et al., 2016; de Sanctis et al., 2016; De Sanctis et al., 2017). Both Himalia and Ceres have 2.7 bands with consistent centers (2.72 ; de Sanctis et al., 2015). Pilorget et al. (2021) also detected this combination of 2.72 and 3.05 bands within the sample of asteroid Ryugu returned by Hayabusa2 mission. The Ryugu sample also contains some trace clasts with unaltered silicates (Brunetto et al., 2023), and the asteroid has been suggested to have formed near to the ice giants before its implantation into the main belt (Hopp et al., 2022; Nesvorný et al., 2024). Figure 6 shows a comparison to 67P (Kurokawa et al., 2020), a presumably non-aqueously-altered object with absorptions between 3.0-3.5 attributed to ammonium-bearing salts (Poch et al., 2020).
@@
The 2.72 features on Ceres and Himalia are similar to some non-ammoniated meteoritic phyllosilicates. The center of the 2.7 feature caused by meteoritic phyllosilicates depends on the Mg/Fe content and therefore the degree of aqueous alteration. Initially anhydrous silicates become Fe-rich serpentine with longer wavelength centers, with continued alteration increasing abundance and replacing Fe- with Mg to form antigorite and magnetite before ultimately forming clays like saponite (Howard et al., 2011). Heavily altered CI chondrites fit farther along this progression compared with typically less altered CM chondrites, though considerable variation exists in the 3 band shapes of these groups of meteorites(Takir et al., 2019). Cere’s feature has been interpreted as similar in strength and location to CI and heavily altered CM chondrites (McSween et al., 2018). Himalia’s 2.7 band has the same center as Ceres’s, suggesting both objects have phyllosilicates with similar Mg/Fe content. The band depth and shape of Himalia’s 2.7 feature, however, may be more similar to the heavily altered CMs and C2-Ungrouped meteorites such as Tagish Lake and Essebi.
@@
Himalia’s 2.7 absorption feature shares some similarities with meteoritic materials, but the 3.05 feature has not been detected from any known meteorite. However, Himalia’s combined 3 band complex is well-matched by laboratory samples of ammoniated phyllosilicates (De Angelis et al., 2021). Figure 7 shows that samples of illite, a clay mineral, can provide a reasonable fit to Himalia’s spectrum from 2.6-2.9 microns. However, beyond 2.9 microns, illite’s narrower absorption feature diverges from Himalia’s wider band complex. When ammoniated in a laboratory setting, illite’s absorption broadens and shows the key 3.05 micron absorption minimum. This feature survives heating up to (De Angelis et al., 2021). We note that the ammonium content on Himalia is likely significantly lower than what is contained in the ammoniated-illite sample, as Himalia’s spectrum is better described by a linear combination of untreated illite with the ammoniated sample (Figure 7.
@@
Other, more complicated mixtures that include ammoniated materials can also fit the 2.7 and 3.05 m bands. We do not consider the case of arbitrarily many absorbers (and thus arbitrary mixture complexity), but we do note that plausible matches are also provided by mixtures of non-ammoniated phyllosilicates with a separate ammonium-bearing material. The presence of such a mixture would imply that the phyllosilicates did not interact with NH-solutions during the parent body’s aqueous processing, a geochemically incoherent situation if both materials were endogenic to the parent body. Yet this two-component example does provide a noteworthy spectral match, as illustrated in Figure 7. A combination of the hydrated C2-Ungrouped chondrite Essebi and an anhydrous ammonium-bearing material (represented by 67P’s nucleus Kurokawa et al., 2020) provides a strong fit to the shape and location of Himalia’s 3.05 feature. The Essebi/67P combination also reproduces the overall width and location of Himalia’s 3 complex, but with two caveats: Essebi contains a strong organics feature at 3.4 and has a longer wavelength band minimum.
@@
We conclude that hydrated silicates and ammonia-bearing materials (either as ammoniated phyllosilicates or as mixed components) provides the best fit to Himalia’s band complex. We do not find any other single material that fits the overall shape of Himalia’s entire band complex from 2.7-3.8 microns. For its simplicity as a single-absorber fit, we favor ammoniated phyllosilicates as the cause of the band complex. While the laboratory samples of illite provide a useful demonstration of the ammoniation process, meteorites display a variety of phyllosilicates in their matrices. Himalia’s surface is likely to be at least similarly complex.
@@ 5 The Himalia Family’s Compositional Heterogeneity
Himalia and its family dominate the mass of the irregular satellite population, and therefore the overall abundance of materials in the system. The discovery of spectral heterogeneity within this family is surprising, particularly given the similarities between Himalia and Elara in the VNIR (Bhatt et al., 2017).
A comparison of the Himalia family members is shown in Figure 8. Himalia, Elara, and Lysithea all have major spectral features that clearly differ from each other. These differences are most easily seen by comparing Himalia and Lysithea, which have different shapes from 2.6-2.9 . Additional differences can be seen near 3 : Himalia’s rounded band is qualitatively different from Lysithea’s sharper, more jagged features in this region. Despite their distinctions, Elara suggests that these two materials are somehow linked. Elara’s absorption feature is directly matched by a simple (50/50) mean of Himalia and Lysithea’s band complexes (Figure 8, Bottom). The Himalia/Lysithea blend describes both Elara’s 2.7 feature and Elara’s intermediate band shape/center from 2.9-3.2 . Elara’s surface composition is, therefore, apparently related to both the larger Himalia and the smaller Lysithea.
One hypothesis is that Lysithea represents material that was exogenic to Himalia’s parent body, possibly sourced from the impactor of the family forming event itself. In that scenario, Elara’s intermediate spectrum would be a mixture of the impactor and the Himalian parent. However, that scenario is inconsistent with Himalia and Elara’s strong spectral similarity at all other wavelengths. Lysithea has a different color from 0.7-1.5 compared with both Elara (Figure 3) and Himalia (Bhatt et al., 2017). If bulk Lysithean material was simply combined with Himalian material, all of Elara’s spectral properties would be intermediate between Himalia and Lysithea. Instead, Elara’s 3 absorber is intermediate between Himalia and Lysithea, but its color and absorption is similar to Himalia. Furthermore, all three objects have the 3.9 feature (which could relate to carbonates). The Himalia family members, therefore, appear compositionally linked and have aqueous alteration products that vary with size.
@@
To explain Elara’s combination of Himalia-like colors and blend of Himalian/Lysithean 3 bands, we hypothesize that these objects were all primarily formed from the same parent body. If Himalia, Elara, and Lysithea all reflect material from the same parent body, our observed progression of hydration features agrees with the hypothesis of Vilas & Hendrix (2024), who proposed that these bodies represent the parent body’s core, mantle, and crust material, respectively.
@@
Future searches for surface heterogeneity and measurements of smaller family members are necessary to constrain family variability: it is yet unknown if there are any consistent spectral types within the Himalia family or a continuum of spectra between two endmembers. Constraining compositional trends would be useful to determine how aqueous alteration occurred on the Himalia parent body. Alteration may have been a function of depth on some irregular satellite parent bodies (e.g., Vilas & Hendrix, 2024; Sharkey et al., 2023) or perhaps occurred in more localized pockets (e.g., like the CV3 Mokoia parent body, Tomeoka & Ohnishi, 2011).
@@
Overall, we propose that the Himalia family parent body was similar to Ceres in terms of containing water, organics, and NH-bearing materials. Furthermore, the location of Himalia’s 2.7 feature is in close agreement with the same feature on Ceres, suggesting both objects experienced similar levels of aqueous alteration. Yet Himalia and Ceres differ in the shape and depth of the 3.05 region. It is possible that the spectral distinctions reflect differences in initial abundances of NH-compounds, but Ceres and Himalia have also likely experienced divergent geological processing. Discerning whether or not Himalia and Ceres could have formed from the same primordial reservoirs provides novel opportunities to explore both objects’ migratory histories and also positions the smaller Himalia family members as possible samples of the interiors of Ceres-like objects.
@@
The ultimate origins of Himalia, and by extension NH in the Jovian system, are still unclear. Beyond the comparisons to Ceres and other inner-solar-system asteroids, we also note similarities between Himalia and the Jovian inner satellite Amalthea from (Takato et al., 2004). Amalthea’s reflectance spectrum has not been characterized between , limiting the constraining power of this comparison considerably. However, the similarities in possible shape and location of Amalthea’s absorption band to Himalia’s are intriguing. Amalthea is thought to represent material that accreted close to Jupiter during the planet’s formation (Greenberg, 2010), and provides a critical constraint on the types of planetesimals that accreted prior to Jupiter’s hypothesized migration. Understanding if Himalia includes any Amalthea-like materials would inform whether or not Himalia itself formed near to Jupiter prior to its capture. Given the potential implications for the Himalia’s provenance, as well as the possible effects of Himalian dust on Gailelan surfaces, future spectroscopic assessment of Amalthea and the other small inner satellites is crucial to understand the Jovian satellite system as a whole.

### v2

@@ 3 Spectral Groupings and Band Identifications
Right: Continuum-divided reflectances using a polynomial of order 2-4 fit on either side of major absorption bands. Objects are normalized to have reflectance equal to 1 at 2.5 microns, and displayed with offsets for ease of comparison. Targets are arranged according to spectral characteristics, not orbital groupings. Major band complexes near 3.0 microns differentiate between the different material categories. The steep red slopes of Carme, Sinope, and Themisto (Group 1) are similar to D-type asteroids or “red” type Jovian Trojans. Ananke (Group 2) shows a broad feature at similar wavelengths to Lysithea and Pasiphae but with a very different, rounded shape. Lysithea and Pasiphae (Group 3) show deep features centered on 3.0 microns similar to less-red Jovian Trojans. Himalia and Elara (Group 4) display a clear 2.7 micron feature, detected previously on aqueously altered inner-solar system materials but never identified in the Jovian system.
@@ 3.1 Continua Definitions
The processed, thermally-corrected reflectance spectrum of each target is shown in Figure 3 (Left). Proper analysis of absorption features requires clear definitions of the spectral continuum. To aid reader comparisons, we define a “global” continuum, used in constructing Figure 3 (Right). This continuum was fit per-object with a low order (2-4) polynomial on either side of the identified 3 micron complex, typically near 2.5 and 4.5 . Dividing the spectrum by this polynomial transforms the spectra from relative reflectance to continuum-divided reflectances. In continuum-divided space, spectral features are identified via strong divergence from unity.
@@
The edges of all other bands were identified by eye, and continuum points were defined as the average within a 0.01 um box on either side of the feature. A linear continuum was fit to these points and divided out, then a Gaussian was fit to identify the center and depth from the local continuum. Due to its asymmetry, we quantify Himalia’s 2.7 feature uniquely. Its center is defined by fitting a Gaussian near its sharp minimum, and its depth is defined relative to the linear continuum of the overall 3 band complex.
@@ 3.2 Spectral Groups
We find that the irregular satellites can be roughly described by four groups according to their spectral slopes and 3 absorption features. These groupings include some that appear only within individual families, and some that appear in multiple families. Band parameters for all features observed on individual satellites are given in Table 2. The first group is represented by Sinope, Carme, and Themisto, which are all red-Trojan-like objects and are all in different orbital families. Ananke comprises the second group, displaying a deeper, rounded 3.0 micron band and relatively neutral spectral slopes. The third group includes Pasiphae, a retrograde satellite, and Lysithea, a prograde satellite, which have 3.0 features centered similarly to Ananke, but with redder spectral slopes and a sharper shape from 2.7-3.3 . Finally, Himalia and Elara, both in the prograde Himalia family, show a unique band complex that includes a previously unobserved sharp 2.7 micron band that transitions into the broad, round feature from 2.8-3.5 microns that was observed by Brown & Rhoden (2014).
Sinope and Carme have similarly steeply increasing spectra from 0.7-2.5 microns, and shallow absorption features near 3.0 and 3.4 , akin to the “red” class of Jupiter Trojans (Emery et al., 2011; Wong et al., 2024). Notably, Themisto, a small () individual prograde satellite interior to the Himalia family, groups closely with these more weakly bound retrograde satellites. The spectral slopes of Sinope and Carme, when overlapping, are consistent with previous measurements at VNIR wavelengths by Sharkey et al. (2023). Carme has a slightly deeper band at 3.0 microns than Sinope and Themisto (depths of vs. ), but they all have similar 3.4 depths ().
@@
| Attribution | Organics | Organics | Organics | Organics | Organics | - | Organics? | - | |
@@
Ananke has the most neutral (least-red) spectral slope within our sample and a deep, rounded 3.0 absorption band. Its 3 band depth () and center () is similar to the redder Pasiphae and Lysithea, though some distinctions in band shape are present. A narrow feature at 2.86 is overprinted within Ananke’s rounded 3 complex. Beyond 3.6 , Ananke’s spectrum is neutral and featureless. Our observations of Ananke differ from the ground-based observations of Sharkey et al. (2023) from 0.7-1.6 microns, which showed a slope that is approximately 10% redder. A similar discrepancy was noted by Wong et al. (2024) when comparing their JWST observations of the Trojan Polymele to Sharkey et al. (2019)’s ground-based observations with the same methodology and setup. Given that both Polymele and Ananke have similar brightnesses V 19, it is possible that those short wavelength slope discrepencies are indicative of systematic errors in the ground-based observations of these very faint objects. Previous study from Marsset et al. (2020) characterized the spectral slope uncertainties for similar ground-based observations to be .
Pasiphae (Pasiphae family) groups closely with Lysithea (Himalia family) in terms of spectral slope and absorption features. They trace closely together shortward of . Pasiphae displays a steeper red slope from 3.5 to 5.2 , while Lysithea rolls over to neutral (flat) reflectance, similar to Ananke. Pasiphae and Lysithea display deep absorption features () with similar positions to Ananke’s.
A comparison of Pasiphae’s and Lysithea’s continuum-divided bands are shown in figure 4, highlighting their similar absorption shapes. Both Pasiphae and Lysithea have additional faint features of 3% overprinted near 3.01 . The cause(s) for the major 3 absorptions on Pasiphae and Lysithea are not clear. Lysithea has an additional feature near 3.9 , which is similar to features caused by carbonates. However, carbonates also display absorptions near 3.4 which we do not detect, confusing a clear identification. Lysithea also shows a narrow feature at 2.63 whose origin is also unclear, although it is near trace features seen in some carbonates (Bishop et al., 2021).
@@
We also find a trace absorption at on Himalia and Elara. This feature is similar to one discovered on the Jupiter Trojan Eurybates at by Wong et al. (2024). Figure 5 compares our observations of Himalia and Elara with the the Wong et al. (2024) observations of Eurybates, as processed via our reduction pipeline. We reprocess this spectrum only for methodological consistency between these datasets but do not note differences between reductions. The 4.26-4.27 features are similar to those caused by solid-state CO2, which is volatile and unlikely to be stable as a crystalline ice on the relatively warm airless surfaces of Jovian satellites or Trojans (Wong et al., 2024). Since CO2 is implausible as pure surface ice, it is likely complexed with more refractory components, as discussed for the origins of 4.25 and 4.27 features on the Galilean moons (e.g., McCord et al., 1998; Villanueva et al., 2023; Cartwright et al., 2024, 2025; Trumbo & Brown, 2023). The discoveries of the features on Eurybates, Himalia, and Elara place complexed CO2 on several compositionally distinct small bodies near Jupiter.
The 3.0 absorber(s) for Carme, Sinope, Themisto, Pasiphae, and Lysithea are consistent with a variety of OH-bearing species, which we discuss in Sections 7 and 8. The absorption at 3.4 on Pasiphae, Ananke, Themisto, Sinope, and Carme is consistent with complex organics. All three objects in the Himalia family (Himalia, Elara, Lysithea) display a feature near 3.9 whose origins are unclear. Carbonates often present dual features near 3.9 and 3.4 , which may be consistent with Elara but not Himalia nor Lysithea. We leave the causes of the Himalia family’s 3.4 and 3.9 features as open questions in Table 2.
@@ 4 Ammoniated Phyllosilicates in the Himalia Family
Some meteorite materials with significant amounts of adsorbed terrestrial water can reproduce the overall width of Himalia’s band complex. However, we note that experimental studies of both meteorites and terrestrial phyllosilicates find that adsorbed water, and the spectral feature it causes, is easily removed when samples are treated under proper vacuum and desiccation procedures designed to mimic asteroid surface conditions (Beck et al., 2010; Takir et al., 2013, 2019). We therefore disfavor fits to Himalia that use phyllosilicates which are not treated to remove terrestrially adsorbed water.
@@
Himalia’s combination of absorptions at 2.7 and 3.05 is reminiscent of Ceres (Kurokawa et al., 2020, Figure 6)) and Hygiea (Rivkin et al., 2025), where the 3.05 feature is associated with the presence of NH4 in the clay mineral structure (e.g., King et al., 1992; Rivkin et al., 2006; Ammannito et al., 2016; de Sanctis et al., 2016; De Sanctis et al., 2017). Himalia, Ceres, and Hygiea have 2.7 bands with consistent centers (2.72 ; de Sanctis et al., 2015; Rivkin et al., 2025, Table 2). Pilorget et al. (2021) also detected this combination of 2.72 and 3.05 bands within the sample of asteroid Ryugu returned by Hayabusa2 mission. The Ryugu sample also contains some trace clasts with unaltered silicates (Brunetto et al., 2023), and the asteroid has been suggested to have formed near to the ice giants before its implantation into the main belt (Hopp et al., 2022; Nesvorný et al., 2024). Figure 6 shows a comparison to 67P (Raponi et al., 2020), a presumably non-aqueously-altered object with absorptions between 3.0-3.5 attributed to ammonium-bearing salts (Poch et al., 2020).
@@
The 2.72 features on Ceres and Himalia are similar to some non-ammoniated meteoritic phyllosilicates. The center of the 2.7 feature caused by meteoritic phyllosilicates depends on the Mg/Fe content and therefore the degree of aqueous alteration (increased alteration corresponding to lower wavelength band centers). Initially anhydrous silicates become Fe-rich serpentines. Continued alteration increases phyllosilicate abundance and replaces Fe- with Mg to form antigorite and magnetite before ultimately forming clays like saponite (Howard et al., 2011). Heavily altered CI chondrites fit farther along this progression compared with typically less altered CM chondrites, though considerable variation exists in the 3 band shapes of these groups of meteorites (Takir et al., 2019). Ceres’s feature is similar in strength and location to CI and heavily altered CM chondrites (McSween et al., 2018). Himalia’s 2.7 band has the same center as Ceres’s, suggesting both objects have phyllosilicates with similar Mg/Fe content. The band depth and shape of Himalia’s 2.7 feature, however, may be more similar to the heavily altered CMs and C2-Ungrouped meteorites such as Tagish Lake and Essebi.
@@
Himalia’s 2.7 absorption feature shares some similarities with meteoritic materials, but the 3.05 feature has not been detected on any known meteorite. Instead, Himalia’s combined 3 band complex is well-matched by laboratory samples of ammoniated phyllosilicates (De Angelis et al., 2021). Figure 7 shows that samples of illite, a clay mineral, can provide a reasonable fit to Himalia’s spectrum from 2.6-2.9 microns. However, beyond 2.9 microns, illite’s narrower absorption feature diverges from Himalia’s wider band complex. When ammoniated in a laboratory setting, illite’s absorption broadens and shows the key 3.05 micron absorption minimum. The presence of the 3.05 feature constrains the thermal history of Himalia’s surface, as De Angelis et al. (2021) found that the absorption feature only survives temperatures up to . Such a temperature limit is consistent with the presence of Himalia’s feature (Jarvis et al., 2000; Vilas & Hendrix, 2024), which disappears in CM2 materials heated C (673 K) (Hiroi et al., 1993) and has been linked to aqueous alteration processes that did not exceed 327 C (600K) (Vilas & Sykes, 1996). We note that the ammonium content on Himalia is likely significantly lower than what is contained in the ammoniated-illite sample, as Himalia’s spectrum is better described by a linear combination of untreated illite with the ammoniated sample (Figure 7).
@@
Other, more complicated mixtures that include ammoniated materials can also fit the 2.7 and 3.05 m bands. We do not consider the case of arbitrarily many absorbers (and thus arbitrary mixture complexity), but we do note that plausible matches are also provided by linear mixtures of non-ammoniated phyllosilicates with a separate ammonium-bearing material. The presence of such a mixture would imply that the phyllosilicates did not interact with NH-solutions during the parent body’s aqueous processing, a geochemically incoherent situation if both materials were endogenic to the parent body. Yet this two-component example does provide a noteworthy spectral match, as illustrated in Figure 7. A combination of the hydrated C2-Ungrouped chondrite Essebi and an anhydrous ammonium-bearing material (represented by 67P’s nucleus Kurokawa et al., 2020) provides a strong fit to the shape and location of Himalia’s 3.05 feature. The Essebi/67P combination also reproduces the overall width and location of Himalia’s 3 complex, but with two caveats: Essebi contains a strong organics feature at 3.4 and has a longer wavelength band minimum.
@@
We conclude that hydrated silicates and ammonia-bearing materials (either as ammoniated phyllosilicates or as mixed components) provides the best fit to Himalia’s band complex. We do not find any other single material that fits the overall shape of Himalia’s entire band complex from 2.7-3.8 microns. For its simplicity as a single-absorber fit, we favor ammoniated phyllosilicates as the cause of the band complex. While the laboratory samples of illite provide a useful demonstration of the ammoniation process, meteorites display a variety of phyllosilicates in their matrices. Himalia’s surface is likely to contain similarly complex phyllosilicate compositions.
@@ 5 The Himalia Family’s Compositional Heterogeneity
. Note that Lysithea displays a sharper band, overprinted by numerous subtle absorptions, while Elara and Himalia have more rounded features near 3 . Bottom: We averaged Himalia’s and Lysithea’s spectra into a new, combined spectrum that closely matches Elara, with the only exception being Elara’s lack of a 2.63 feature. The 3 band depth of the averaged Himalia/Lysithea absorption was multiplied by a factor of 1.07 to make comparison to the shape of Elara’s spectral features easier.
The Himalia family dominates the mass of Jupiter’s irregular satellites by a factor of (assuming spherical volumes with equal densities for all satellites). The Himalia family therefore stands as a major possible source for carbonaceous material within the Jovian system. But the bulk abundance of material within the Himalia family is unclear, as all three family members have different band complexes. If Himalia, Elara, and Lysithea formed from the same parent body, that would require that their parent experienced heterogeneous aqeuous alteration. Alternatively, their spectral variation may represent a mixture of materials from multiple parent bodies. Here, we explore whether the observed spectral properties of Himalia, Elara, and Lysithea favor a single-parent or multi-parent origin.
A comparison of the Himalia family members’ continuum-removed region is shown in Figure 8. Lysithea presents a single-minimum band, while Himalia and Elara display composite features with minima near 2.7 and 3.0 . The Himalia family’s band properties do not just vary in depth: Himalia has a rounded band that is centered longer than Lysithea’s sharper feature (see Table 2. Elara’s absorption feature is similar to a simple (50/50) mean of Himalia and Lysithea’s band complexes (Figure 8, Right). The Himalia/Lysithea blend describes both Elara’s 2.7 feature and Elara’s intermediate band shape/center from 2.9-3.2 . Elara’s surface appears to contain both Lysithea’s unknown 3.0 absorber and the ammoniated phyllosilicates present on Himalia.
While Elara’s band complex is intermediate between Himalia and Lysithea, its overall spectrum is not. Elara’s spectral slope matches Himalia, while Lysithea diverges from 0.7-1.5 compared with both Elara (Figure 3) and Himalia (Bhatt et al., 2017). The presence of trace absorption features add further complexity: Himalia and Elara display 4.27 features; Himalia, Elara, and Pasiphae all display 3.9 features; Himalia and Lysithea display 2.63 features. The Himalia family members present a complex compositional picture whose only clear spectral trend is the correlation between the presence of ammoniated materials and satellite size.
Elara is not a simple mixture of Lysithean and Himalian materials, but it does show that there is some relationship between their 3 absorbers. We suggest that Elara’s combination of Himalia-like colors and its blend of Himalian/Lysithean 3 bands is best understood as a distinct processing state of material from the Himalia parent body. This interpretation favors the hypothesis that all three objects were primarily formed from the same parent body, with the three objects representing distinct aqueous alteration states. The origins of trace absorption bands remain unclear, but could relate to the specifics of parent body alteration processes.
@@
Regardless of specific origins, the Himalia family contains at least 2 distinct types of OH-bearing materials (Himalia-like and Lysithea-Like) that do not correlate with optical or near-infrared color. If Himalia, Elara, and Lysithea all reflect material from the same parent body, our observed progression of hydration features agrees with the observations and hypothesis of Vilas & Hendrix (2024), who proposed that these bodies represent the parent body’s core, mantle, and crust material, respectively.
@@
Future searches for surface heterogeneity and measurements of smaller family members are necessary to constrain family variability. It is yet unknown if there are any consistent spectral types within the Himalia family or if there is instead a continuous array of compositions between separate endmembers. Constraining compositional trends would be useful to determine how aqueous alteration occurred on the Himalia parent body. Alteration may have occurred as a function of depth on some irregular satellite parent bodies (e.g., Vilas & Hendrix, 2024; Sharkey et al., 2023) or perhaps in more localized pockets (e.g., like the CV3 Mokoia parent body, Tomeoka & Ohnishi, 2011).
@@
We propose that the Himalia family parent body was similar to Ceres in terms of containing water, organics, and NH-bearing materials. Furthermore, the location of Himalia’s 2.7 feature is in close agreement with the same feature on Ceres, suggesting both objects experienced similar aqueous alteration processes. Yet Himalia and Ceres differ in the shape and depth of the 3.05 region. It is possible that the spectral distinctions reflect differences in initial abundances of NH-compounds, but Ceres and Himalia have also likely experienced divergent geological processing. Discerning whether or not Himalia and Ceres could have formed from the same primordial reservoirs constrains both objects’ migratory histories and could position the smaller Himalia family members as possible samples of less-processed Ceres-like objects.
@@
The ultimate origins of Himalia, and by extension NH in the Jovian system, are still unclear. Beyond the comparisons to Ceres and other inner-solar-system asteroids, we also note similarities between Himalia and the Jovian inner satellite Amalthea from (Takato et al., 2004). Amalthea’s reflectance spectrum has not been characterized between , limiting the constraining power of this comparison considerably. However, the similarities in possible shape and location of Amalthea’s absorption band to Himalia’s are intriguing. Amalthea has a low-eccentricity and low-inclination orbit interior to the Galilean satellites, and it is thought to represent material that accreted close to Jupiter during the planet’s formation (Greenberg, 2010). Understanding if Himalia includes any Amalthea-like materials, which may have accreted prior to Jupiter’s hypothesized migration, would inform whether or not Himalia itself formed near to Jupiter prior to its capture. Given the potential implications for Himalia’s provenance, as well as the possible effects of Himalian dust on Galilean surfaces, future spectroscopic assessment of Amalthea and the other small inner satellites is crucial to understand the Jovian satellite system as a whole.